OpenStudy (anonymous):

.'.'.' i will give medal '.'.'. :) The fuel tank of Mr Ong's car is 45 liters. He fills up the fuel tank in Singapore before heading to Penang. The fuel consumption of his car is approximatey 10.8km/liter, find the amount of fuel left (in liters) in the fuel tank after he has travelled (a) 48km (b) 245km Express the amount of fuel (y liters) that was left in the fuel tank after he has travelled x km. If x=-25, what is the value of y ? What will the value of y be when x=558 ? Can you explain the significance of the values you get from the above?

OpenStudy (sheraz12345):

U can use ratio method i guess 1 litre : 10.8km x litre : 48 km

OpenStudy (anonymous):

for a) 48/10.8 =4.4444L, the amount of full left would be 40.5556L for b) 245/10.8= 22.685L, the mount of full left would be 22.315L

OpenStudy (anonymous):

ok i'll try both

OpenStudy (anonymous):

for x= -25km, even though it is negative it has travelled a distance of 25km, therefore y= 45 -(- 25/10.8 = 47.314L which is impossible because mr. ong's car can hold only 45L of fuel for x= 558km, 558/10.8= 51.6666L which is impossible too because mr. ong's car can hold only 45L of fuel

OpenStudy (anonymous):

ok i got it thanks